Please Help!
What would be the determinant of Dx, if Cramer's Rule is used to solve:
3x-2y=5
2x+5y=-3
Thank you for your time!
: Please Help!
What would be the determinant of Dx, if Cramer's Rule is used to solve:
3x-2y=5
2x+5y=-3
Thank you for your time!

Answer by Alan3354(1217)

What would be the determinant of Dx, if Cramer's Rule is used to solve:
3x-2y=5
2x+5y=-3
------
The det = 3*5 - 2*(-2)
= 15 + 4
=19.
Whether or not Cramer's Rule is used is irrelevant, the Det is still 19.